# Latent Adaptive Planner for Dynamic Manipulation

### v1

@@ Acknowledgments
Y. W. was partially supported by NSF DMS-2015577, NSF DMS-2415226, and a gift fund from Amazon. We gratefully acknowledge the support of Lambda, Inc. for providing compute resources for this project.
@@ Appendix A Model Predictive Control (MPC)
$x_{k}=\begin{bmatrix}q_{k}\\ \dot{q}_{k}\end{bmatrix},$
where $q_{k}$ and $\dot{q}_{k}$ represent the joint angles and joint velocities, respectively. Let $u_{k}$ be the control input (torque). Over a prediction horizon of length $N_{p}$, the following cost function is minimized:
$\sum_{k=i}^{i+N_{p}}\Big{[}(\hat{q}_{k}-q_{k})^{T}\,Q_{q}\,(\hat{q}_{k}-q_{k})\;+\;(\hat{\dot{q}}_{k}-\dot{q}_{k})^{T}\,Q_{\dot{q}}\,(\hat{\dot{q}}_{k}-\dot{q}_{k})\;+\;u_{k}^{T}\,Q_{u}\,u_{k}\Big{]},$
@@
$x_{k+1}=A_{k}\,x_{k}+B_{k}\,u_{k}+r_{k},$
where $A_{k}$, $B_{k}$, and $r_{k}$ approximate the robot’s dynamics around a nominal operating point. The following constraints are imposed to ensure feasibility and safety:
$u_{\min}$ $\leq u_{k}\leq u_{\max},$
$q_{\min}$ $\leq q_{k}\leq q_{\max},$
$\dot{q}_{\min}$ $\leq\dot{q}_{k}\leq\dot{q}_{\max},$
$\ddot{q}_{\min}$ $\leq\ddot{q}_{k}\leq\ddot{q}_{\max},$
$q_{0}$ $=q_{\text{current}},\quad\dot{q}_{0}=\dot{q}_{\text{current}},$
where $q_{\min},q_{\max},\dot{q}_{\min},\dot{q}_{\max}$, and so forth denote predefined joint and actuator limits. Collecting the state and input variables into a decision vector
$z_{k}=\begin{bmatrix}x_{k}\\ u_{k}\end{bmatrix},$
@@
$\min_{\{z_{k}\}_{k=i}^{i+N_{p}}}$ $\sum_{k=i}^{i+N_{p}}\bigl{(}z_{k}^{T}W_{k}\,z_{k}+2\,w_{k}^{T}\,z_{k}\bigr{)},$
subject to $x_{k+1}=A_{k}\,x_{k}+B_{k}\,u_{k}+r_{k},$
$u_{\min}\leq u_{k}\leq u_{\max},$
$q_{\min}\leq q_{k}\leq q_{\max},$
$\dot{q}_{\min}\leq\dot{q}_{k}\leq\dot{q}_{\max},$
$\ddot{q}_{\min}\leq\ddot{q}_{k}\leq\ddot{q}_{\max},$
$q_{0}=q_{\text{current}},\quad\dot{q}_{0}=\dot{q}_{\text{current}}.$
The matrices $W_{k}$ and vectors $w_{k}$ encapsulate the quadratic and linear terms derived from expanding the cost function and incorporating the linearized dynamics. Solving this QP at each time step allows the robot to effectively track the learning policy’s references $\hat{q}_{k},\hat{\dot{q}}_{k}$ while honoring physical constraints and preserving safe operation.
@@ B.1 Actuators
Dimensions (mm) $63.5\times 62\times 37$ $75\times 67\times 37.5$ $113\times 113\times 49.7$ $113\times 113\times 49.7$
@@ B.2 Robot A
Body Yaw $z$-axis rotation (0.0, 0.0, 1.1276) $[-1.57,1.57]$
Shoulder Yaw $z$-axis rotation (0.0, $\pm$0.21, 0.117) $[-2.27,2.27]$
Shoulder Pitch $y$-axis rotation (0.0, $\pm$0.15, 0.0) $[-1.57,1.57]$
Elbow Pitch $y$-axis rotation (0.0, 0.4, 0.0) $[-1.57,2.53]$
Wrist Pitch $y$-axis rotation (0.0, 0.375, 0.0) $[-2.44,2.44]$
Wrist Roll $x$-axis rotation (0.01925, 0.0, 0.0) $[-1.57,1.57]$
@@ B.3 Robot B
Body Yaw $z$-axis rotation (0.0, 0.0, 1.1327) $[-1.57,1.57]$
Shoulder Pitch $y$-axis rotation (0.0, $\pm$0.19475, 0.0) $[-1.57,1.57]$
@@
Elbow Pitch $y$-axis rotation (0.02, 0.0, -0.27585) $[-0.3,2.9]$
Wrist Pitch $y$-axis rotation (0.23226, 0.019701, 0.0) $[-1.15,2.2]$
@@ Appendix E Latent Adaptive Planner Algorithms
E.1 Classical Variational Learning of LAP
@@
Algorithm 1 Classical Variational Bayes Learning for LAP
@@ E.2 Variational Replanning Algorithm
Algorithm 2 Variational Replanning for LAP

### v2

@@ Acknowledgments
Y. W. was partially supported by NSF DMS-2015577, NSF DMS-2415226, and a gift fund from Amazon. We gratefully acknowledge Lambda, Inc. for providing compute resources for this project. We thank Sirui Xie for early collaboration and conceptual discussions.
@@ Appendix A Model Predictive Control (MPC)
$\mathbf{x}_{k}=\begin{bmatrix}\mathbf{q}_{k}\\ \dot{\mathbf{q}}_{k}\end{bmatrix}$,
where $\mathbf{q}_{k}$ and $\dot{\mathbf{q}}_{k}$ represent the joint angles and joint velocities, respectively. Let $\mathbf{u}_{k}$ be the control input (torque). Over a prediction horizon of length $N_{p}$, the following cost function is minimized:
$\sum_{k=i}^{i+N_{p}}\Big{[}(\hat{\mathbf{q}}_{k}-\mathbf{q}_{k})^{T}\,Q_{q}\,(\hat{\mathbf{q}}_{k}-\mathbf{q}_{k})\;+\;(\hat{\dot{\mathbf{q}}}_{k}-\dot{\mathbf{q}}_{k})^{T}\,Q_{\dot{q}}\,(\hat{\dot{\mathbf{q}}}_{k}-\dot{\mathbf{q}}_{k})\;+\;\mathbf{u}_{k}^{T}\,Q_{u}\,\mathbf{u}_{k}\Big{]},$
@@
$\mathbf{x}_{k+1}=A_{k}\,\mathbf{x}_{k}+B_{k}\,\mathbf{u}_{k}+\mathbf{r}_{k},$
where $A_{k}$, $B_{k}$, and $\mathbf{r}_{k}$ approximate the robot’s dynamics around a nominal operating point. The following constraints are imposed to ensure feasibility and safety:
$\mathbf{u}_{\min}$ $\leq\mathbf{u}_{k}\leq\mathbf{u}_{\max},$
$\mathbf{q}_{\min}$ $\leq\mathbf{q}_{k}\leq\mathbf{q}_{\max},$
$\dot{\mathbf{q}}_{\min}$ $\leq\dot{\mathbf{q}}_{k}\leq\dot{\mathbf{q}}_{\max},$
$\ddot{\mathbf{q}}_{\min}$ $\leq\ddot{\mathbf{q}}_{k}\leq\ddot{\mathbf{q}}_{\max},$
$\mathbf{q}_{0}$ $=\mathbf{q}_{\text{current}},\quad\dot{\mathbf{q}}_{0}=\dot{\mathbf{q}}_{\text{current}},$
where $\mathbf{q}_{\min},\mathbf{q}_{\max},\dot{\mathbf{q}}_{\min},\dot{\mathbf{q}}_{\max}$, and so forth denote predefined joint and actuator limits. Collecting the state and input variables into a decision vector
$\mathbf{z}_{k}=\begin{bmatrix}\mathbf{x}_{k}\\ \mathbf{u}_{k}\end{bmatrix}$,
@@
$\min_{\{\mathbf{z}_{k}\}_{k=i}^{i+N_{p}}}$ $\sum_{k=i}^{i+N_{p}}\bigl{(}\mathbf{z}_{k}^{T}W_{k}\,\mathbf{z}_{k}+2\,\mathbf{w}_{k}^{T}\,\mathbf{z}_{k}\bigr{)},$
subject to $\mathbf{x}_{k+1}=A_{k}\,\mathbf{x}_{k}+B_{k}\,\mathbf{u}_{k}+\mathbf{r}_{k},$
$\mathbf{u}_{\min}\leq\mathbf{u}_{k}\leq\mathbf{u}_{\max},$
$\mathbf{q}_{\min}\leq\mathbf{q}_{k}\leq\mathbf{q}_{\max},$
$\dot{\mathbf{q}}_{\min}\leq\dot{\mathbf{q}}_{k}\leq\dot{\mathbf{q}}_{\max},$
$\ddot{\mathbf{q}}_{\min}\leq\ddot{\mathbf{q}}_{k}\leq\ddot{\mathbf{q}}_{\max},$
$\mathbf{q}_{0}=\mathbf{q}_{\text{current}},\quad\dot{\mathbf{q}}_{0}=\dot{\mathbf{q}}_{\text{current}}.$
The matrices $W_{k}$ and vectors $\mathbf{w}_{k}$ encapsulate the quadratic and linear terms derived from expanding the cost function and incorporating the linearized dynamics. Solving this QP at each time step allows the robot to effectively track the learning policy’s references $\hat{\mathbf{q}}_{k},\hat{\dot{\mathbf{q}}}_{k}$ while honoring physical constraints and preserving safe operation.
@@ B.1 Actuators
Dimensions (mm) $63.5\times 62\times 37$ $75\times 67\times 37.5$ $\times 113\times 49.7 113×113$$\times 113\times 49.7$
@@ B.2 Robot A
Body Yaw $z$-axis rotation (0.0, 0.0, 1.1276) [1.57,1.57][-1.57,1.57]
Shoulder Yaw $z$-axis rotation (0.0, $\pm$0.21, 0.117) [2.27,2.27][-2.27,2.27]
Shoulder Pitch $y$-axis rotation (0.0, $\pm$0.15, 0.0) [1.57,1.57][-1.57,1.57]
Elbow Pitch $y$-axis rotation (0.0, 0.4, 0.0) [1.57,2.53][-1.57,2.53]
Wrist Pitch $y$-axis rotation (0.0, 0.375, 0.0) [2.44,2.44][-2.44,2.44]
Wrist Roll $x$-axis rotation (0.01925, 0.0, 0.0) [1.57,1.57][-1.57,1.57]
@@ B.3 Robot B
Body Yaw $z$-axis rotation (0.0, 0.0, 1.1327) [1.57,1.57][-1.57,1.57]
Shoulder Pitch $y$-axis rotation (0.0, $\pm$0.19475, 0.0) [1.57,1.57][-1.57,1.57]
@@
Elbow Pitch $y$-axis rotation (0.02, 0.0, -0.27585) [0.3,2.9][-0.3,2.9]
Wrist Pitch $y$-axis rotation (0.23226, 0.019701, 0.0) [1.15,2.2][-1.15,2.2]
Figure 5: Coordinate system of Robot B
@@ Appendix E Latent Adaptive Planner Algorithms
E.1 Classical Variational Bayes Learning of LAP
@@
Algorithm 1 LAP: Classical Variational Bayes Learning
@@ E.2 Variational Replanning Algorithm
Algorithm 2 LAP: Variational Replanning